Jeff Corporation purchased a patent for $375,000 on May 1, 2019. It has a useful life of 10 years. What is the balance in the patent account at December 31, 2021? (This is their only patent.)
Answer #1

Total amortization expense should be calculated first.

Purchase price = $375,000

Useful life = L = 10 years

Months in a year = M = 12

Months between May 1, 2019 and Dec 31, 2021 = B = 32

By the use of formula as below:

Total amortization expense = (Purchase price ÷ L) × (B / M)

                                                = (375,000 ÷ 10) × (32 / 12)

                                                = 37,500 × (32 / 12)

                                                = (37,500 × 32) / 12

                                                = 1,200,000 / 12

                                                = 100,000

This is to be subtracted from the purchase price in order to get the balance amount.

Balance amount = Purchase price – Total amortization expense

                           = 375,000 – 100,000

                           = $275,000 (Answer)